# A Strong View-Free Baseline Approach for Single-View Image Guided Point Cloud Completion

### v1

@@ I Introduction
Background. Recent advancements in 3D perception have significantly increased research and industrial interest in point cloud completion, a crucial task for robotics, autonomous driving, and augmented reality [1][2]. However, real-world sensor data often contains incomplete point clouds due to factors such as occlusions, restricted viewpoints, and noise [3]. To address these issues, deep learning–based methods have substantially enhanced reconstruction quality by understanding complex spatial relationships and learning how to map sparse inputs to dense 3D shapes [4][5], paving the way for more advanced network architectures that achieve state-of-the-art performance in point cloud completion.
@@
Existing research on deep learning for point cloud completion has predominantly utilized an encoder-decoder architecture, where the encoder transforms incomplete input into a latent feature space that captures both global structure and local geometric details, while the decoder reconstructs a dense 3D representation from these features [6]. Building on this foundation, methods such as AtlasNet [7], FoldingNet [8], and TopNet [9] have introduced alternative decoder architectures designed to enhance the representation of surface geometries, leading to sharper and more coherent reconstructions. More recently, several studies have explored complementary techniques, including attention mechanisms, to refine the modeling of point relationships and improve inference performance [10]. However, the quality of the latent representation in the encoder-decoder framework, as commonly found in existing works, can be a limiting factor in enhancing generation performance. Constructing a latent representation with rich information, while often non-trivial, is crucial for effective point cloud completion.
@@
To construct a latent representation with richer information, some researches adopt multi-modal design to fuse inputs from multiple sources, of which a popular design is view-based methods. A popular approach within this design is view-based methods, which incorporate additional sources of information to enhance point cloud representations. Contrary to traditional deep learning methods which only take incomplete point clouds as inputs, view-based point cloud completion methods, typically require images as additional input alongside the incomplete point cloud. Single-view images provide complementary geometric cues, helping the model to infer missing structures that may be unclear from partial 3D data alone. These methods leverage image features to address occlusions, maintain viewpoint consistency, and often improve completion accuracy by guiding the network toward a more informed reconstruction. For instance, ViPC [11] generates a coarse point cloud from a single-view image before refinement and employs vector concatenation to fuse the two modalities. DMF-Net [12] adopts a dual-channel approach, with one channel processing the partial point cloud and the other processing an image of the object, placing greater emphasis on fusion balance through a symmetric design. CSDN [13] utilizes a shape-transfer mechanism to extract shape characteristics from the input image, guiding the completion process. The image used in CSDN input, which contains the missing parts of the point cloud, assists the network in identifying the regions requiring completion. Song et al. [14] leverages a pretrained CLIP [15] model in a multimodal fusion setting to incorporate both visual and textual information for guiding point cloud completion. In these methods, images, serving as 2D projections of 3D objects, provide additional information about the objects’ appearance and geometry to facilitate the completion process.
@@
Challenges in SVIPC. While incorporating images alone may enhance the completion process by providing additional information, it also introduces additional concerns and limitations. For instance, the performance of view-based methods heavily depends on the quality of input images and the chosen viewpoint, and blurry, occluded, or unfavorably angled images can negatively impact performance [16]. Additional efforts are required for camera calibration, as misalignment or calibration errors can degrade completion performance [17]. Subsequently, images alone may not provide sufficient geometrical information in certain scenarios, especially for 3D shapes lacking texture or containing ambiguous regions [18]. The process of capturing and processing image inputs introduces significant computational complexity, which raises concerns regarding the scalability of SVIPC in real-world applications. These challenges often become even more pronounced, where obtaining a high-quality image specifically associated with the partial point cloud may prove difficult or even infeasible. Surprisingly, there are few studies in the literature that explore the necessity of single-view image guidance in point cloud completion tasks, which significantly motivates our work.
@@
Our Approach & Contributions. To explore this necessity, we propose a strong view-free baseline method for SVIPC. We introduce a multi-branch 3D encoder network that operates solely on partial point clouds. The 3D encoder, built upon PointNet++ [19], extracts multi-scale geometric features and captures both fine-grained local details and global structures. These features are then progressively refined through self-fusion, which leverages multi-stage attentions to guide the completion process by fusing information within the partial point clouds. We term this approach “self-fusion,” as it enriches feature representations by integrating geometric cues directly from the input data. This hierarchical fusion mechanism ensures that the model dynamically adjusts its focus, strengthening representations at different abstraction levels. The multi-branch design enables diverse feature learning, where different stages specialize in capturing different geometric aspects. By integrating these complementary representations, our approach improves reconstruction robustness without relying on external view cues.
@@ II Related Work
Point cloud completion has been extensively studied in 3D computer vision due to its significance in applications such as autonomous driving, robotics, and scene reconstruction [1, 3]. Early approaches to shape completion relied on explicit geometric descriptors or shape retrieval from large datasets, with methods such as data-driven structural priors [20] and primitive shape-based reconstructions [21], though they often struggled with generalizing to complex and diverse geometries. With the advent of deep learning, encoder-decoder architectures such as PCN [22] have become prominent, as they eliminate assumptions about structural priors. Researchers have since proposed a variety of architectures and loss functions [23, 24, 25] to further optimize point cloud completion performance. Subsequent works, including AtlasNet [7] and TopNet [26], advanced point cloud generation by leveraging parametric surface elements and hierarchical structures.
Transformer-based methods have also gained traction. For example, models like PoinTr [10] employ geometry-aware transformers to capture local relationships in 3D spaces, while SnowflakeNet [27] refines point distributions through iterative splitting to produce highly detailed reconstructions. Recent advancements such as VRCNet [28] incorporate dual-path architectures for probabilistic modeling, enhancing both accuracy and robustness.
The integration of auxiliary 2D image data for point cloud completion represents a new direction. Initiated by ViPC [11], view-guided completion introduced the use of single-view reconstructions as priors. However, its reliance on an intermediate coarse point cloud generation can limit performance. Building on this foundation, subsequent works have explored more advanced fusion strategies; for instance, XMFNet [29] leverages a Dynamic Graph CNN for point cloud feature extraction and integrates multiple cross and self-attention layers to achieve single-level feature fusion between image and point cloud modalities, establishing a strong baseline for multimodal 3D reconstruction. More recently, additional view-guided methods have emerged. For example, CSDN [13] leverages shape-transfer to propagate geometric details from global shape to point cloud domain and dual-refinement strategies to iteratively enhance global shape structures and local details; Fine-grained CLIP [14] exploits textual cues in combination with images for more contextually aware shape completion; DMF-Net [30] and EGIInet [31] introduce dual-channel modality fusion and shape-aware upsampling to ensure finer geometric consistency for further performance improvement.
@@ III Our Approach
3D Encoder. Taking an incomplete point cloud as input, the initial features are extracted using a 3D encoder. Our 3D encoder follows a modular design, allowing the use of various feature extractors such as PointNet++ [19] and DGCNN [4]. Inspired by SnowflakeNet, our 3D encoder adopts a Set Abstraction (SA) layer with K-NN. Unlike SnowflakeNet, which retains only the final feature, our encoder retains both intermediate feature representations and the final encoded feature, as shown in the orange and green boxes in Figure 1. In contrast to the single-feature extraction in SnowflakeNet and XMFNet, our 3D encoder extracts hierarchical features at multiple levels, capturing both global structural information and local fine-grained details from the input point cloud. In the two-branch configuration, each branch processes the input point cloud independently through the 3D encoder, ensuring diverse feature extraction. These multi-scale representations are then grouped by hierarchical level before being processed in the self-fusion network.
@@
Point Cloud Generation Decoder. The decoder reconstructs the complete point cloud by estimating the missing points and integrating them with the input partial cloud. Since decoding the acquired fusion features requires flexibility and a strong learning capacity, we adopt a decoder architecture similar to XMFNet [32], designed to process features and learn their implicit shape representations to predict 3D coordinates. Our plug-and-play attention-enhanced self-fusion network is compatible with various other point cloud generation decoders, such as one used by PointAttN [32].
@@ IV-A Datasets
All experiments are conducted on the ShapeNet-ViPC dataset, which contains 38,328 objects from 13 categories. Each object is associated with 24 partial point clouds and corresponding images, generated under 24 viewpoints, following the previous works’ standard setting [11, 13, 29].
@@ IV-B Evaluation Metrics
Following the literature, the performance is evaluated using Chamfer Distance (CD) and F-Score. CD measures the average squared distance between points in the predicted and ground truth point clouds, assessing reconstruction accuracy. F-Score computes overlap precision and recall at a fixed threshold, reflecting structural similarity [11, 31].
@@
TABLE II: Mean F-Score @ 0.001$\uparrow$ on ShapeNet-ViPC
@@ IV-C Performance
Table I and Table II summarize our quantitative comparison results for ShapeNet-ViPC dataset, where we report the best officially reported performance of each method based on prior work [31]. In these comparisons, we evaluate our method alongside existing view-guided point cloud completion approaches, including ViPC [11], CSDN [13], XMFNet [29] and EGIInet [31]. Additionally, we benchmark against general point cloud completion methods, such as AtlasNet [33], FoldingNet [34], PCN [6], TopNet [35], PF-Net [36], MSN [37], GRNet [38], PoinTr [39], PointAttN [32], SDT [40] and Seedformer [41]. Our solution consistently outperforms methods mentioned above in terms of both CD and F-score, demonstrating strong reconstruction accuracy across multiple categories.
@@ IV-E Ablation Study
To assess the impact of different point feature feature extractors and generation decoders on this task, we conduct ablation studies using different combinations of feature extractors (DGCNN [4] and PointNet++ [19]) and generation decoders (XMFNet-based and Transformer-based [32]). As shown in Table IV, the self-fusion network demonstrates strong compatibility with different feature encoders and point cloud generation decoders, highlighting its flexibility and robustness. This study further validates the generality of the proposed self-fusion network across different architectural configurations.
@@
As extensively studied, vanilla CD is vulnerable to outliers, leading to the development of several CD variants including DCD [42], HyperCD [25], InfoCD [24] and LandauCD [23]. To assess the impact of these variants as loss functions, we conduct an ablation study using a two-branch configuration, with results summarized in Table V. Our findings suggest that, without hyperparameter tuning, these CD variants provide only marginal improvements. Given the substantial effort required for hyperparameter optimization and our primary focus on pipeline design, we adopt vanilla CD as the default loss function for simplicity and efficiency.

### v2

@@ I Introduction
Background. Recent advancements in 3D perception have significantly increased research and industrial interest in point cloud completion, a crucial task for robotics, autonomous driving, and augmented reality [xu2022fpcc][hou2025mobile]. However, real-world sensor data often contains incomplete point clouds due to factors such as occlusions, restricted viewpoints, and noise [lin2022cosmos]. To address these issues, deep learning–based methods have substantially enhanced reconstruction quality by understanding complex spatial relationships and learning how to map sparse inputs to dense 3D shapes [wang2019dynamic][zhang2021point], paving the way for more advanced network architectures that achieve state-of-the-art performance in point cloud completion.
@@
Existing research on deep learning for point cloud completion has predominantly utilized an encoder-decoder architecture, where the encoder transforms incomplete input into a latent feature space that captures both global structure and local geometric details, while the decoder reconstructs a dense 3D representation from these features [yuan2018pcn]. Building on this foundation, methods such as AtlasNet [AtlasNet], FoldingNet [yang2018foldingnet], and TopNet [tchapmi2019topnet] have introduced alternative decoder architectures designed to enhance the representation of surface geometries, leading to sharper and more coherent reconstructions. More recently, several studies have explored complementary techniques, including attention mechanisms, to refine the modeling of point relationships and improve inference performance [PoinTr]. However, the quality of the latent representation in the encoder-decoder framework, as commonly found in existing works, can be a limiting factor in enhancing generation performance. Constructing a latent representation with rich information, while often non-trivial, is crucial for effective point cloud completion.
@@
To construct a latent representation with richer information, some researches adopt multi-modal design to fuse inputs from multiple sources, of which a popular design is view-based methods. A popular approach within this design is view-based methods, which incorporate additional sources of information to enhance point cloud representations. Contrary to traditional deep learning methods which only take incomplete point clouds as inputs, view-based point cloud completion methods, typically require images as additional input alongside the incomplete point cloud. Single-view images provide complementary geometric cues, helping the model to infer missing structures that may be unclear from partial 3D data alone. These methods leverage image features to address occlusions, maintain viewpoint consistency, and often improve completion accuracy by guiding the network toward a more informed reconstruction. For instance, ViPC [zhang2021view] generates a coarse point cloud from a single-view image before refinement and employs vector concatenation to fuse the two modalities. DMF-Net [lu2024mmcnet] adopts a dual-channel approach, with one channel processing the partial point cloud and the other processing an image of the object, placing greater emphasis on fusion balance through a symmetric design. CSDN [zhu2023csdn] utilizes a shape-transfer mechanism to extract shape characteristics from the input image, guiding the completion process. The image used in CSDN input, which contains the missing parts of the point cloud, assists the network in identifying the regions requiring completion. Song et al. [song2023fine] leverages a pretrained CLIP [radford2021learning] model in a multimodal fusion setting to incorporate both visual and textual information for guiding point cloud completion. In these methods, images, serving as 2D projections of 3D objects, provide additional information about the objects’ appearance and geometry to facilitate the completion process.
@@
Challenges in SVIPC. While incorporating images alone may enhance the completion process by providing additional information, it also introduces additional concerns and limitations. For instance, the performance of view-based methods heavily depends on the quality of input images and the chosen viewpoint, and blurry, occluded, or unfavorably angled images can negatively impact performance [fu2023vapcnet]. Additional efforts are required for camera calibration, as misalignment or calibration errors can degrade completion performance [tang2025calibration]. Subsequently, images alone may not provide sufficient geometrical information in certain scenarios, especially for 3D shapes lacking texture or containing ambiguous regions [zhou2024position]. The process of capturing and processing image inputs introduces significant computational complexity, which raises concerns regarding the scalability of SVIPC in real-world applications. These challenges often become even more pronounced, where obtaining a high-quality image specifically associated with the partial point cloud may prove difficult or even infeasible. Surprisingly, there are few studies in the literature that explore the necessity of single-view image guidance in point cloud completion tasks, which significantly motivates our work.
@@
Our Approach & Contributions. To explore this necessity, we propose a strong view-free baseline method for SVIPC. We introduce a multi-branch 3D encoder network that operates solely on partial point clouds. The 3D encoder, built upon PointNet++ [qi2017pointnet++], extracts multi-scale geometric features and captures both fine-grained local details and global structures. These features are then progressively refined through self-fusion, which leverages multi-stage attentions to guide the completion process by fusing information within the partial point clouds. We term this approach “self-fusion,” as it enriches feature representations by integrating geometric cues directly from the input data. This hierarchical fusion mechanism ensures that the model dynamically adjusts its focus, strengthening representations at different abstraction levels. The multi-branch design enables diverse feature learning, where different stages specialize in capturing different geometric aspects. By integrating these complementary representations, our approach improves reconstruction robustness without relying on external view cues.
@@ II Related Work
Point cloud completion has been extensively studied in 3D computer vision due to its significance in applications such as autonomous driving, robotics, and scene reconstruction [xu2022fpcc, lin2022cosmos, zhang2024deep, gao2025safecoop]. Early approaches to shape completion relied on explicit geometric descriptors or shape retrieval from large datasets, with methods such as data-driven structural priors [Data-Driven, yue2024understanding] and primitive shape-based reconstructions [veryold], though they often struggled with generalizing to complex and diverse geometries. With the advent of deep learning, encoder-decoder architectures such as PCN [PCN] have become prominent, as they eliminate assumptions about structural priors. Researchers have since proposed a variety of architectures and loss functions [lin2024loss, lin2023infocd, lin2023hyperbolic, lin2024hyperbolic, zhang2025gps] to further optimize point cloud completion performance. Subsequent works, including AtlasNet [AtlasNet] and TopNet [TopNet], advanced point cloud generation by leveraging parametric surface elements and hierarchical structures.
Transformer-based methods have also gained traction. For example, models like PoinTr [PoinTr] employ geometry-aware transformers to capture local relationships in 3D spaces, while SnowflakeNet [SnowflakeNet] refines point distributions through iterative splitting to produce highly detailed reconstructions. Recent advancements such as VRCNet [Variational] incorporate dual-path architectures for probabilistic modeling, enhancing both accuracy and robustness.
The integration of auxiliary 2D image data for point cloud completion represents a new direction. Initiated by ViPC [zhang2021view], view-guided completion introduced the use of single-view reconstructions as priors. However, its reliance on an intermediate coarse point cloud generation can limit performance. Building on this foundation, subsequent works have explored more advanced fusion strategies; for instance, XMFNet [XMFNet] leverages a Dynamic Graph CNN for point cloud feature extraction and integrates multiple cross and self-attention layers to achieve single-level feature fusion between image and point cloud modalities, establishing a strong baseline for multimodal 3D reconstruction. More recently, additional view-guided methods have emerged. For example, CSDN [zhu2023csdn] leverages shape-transfer to propagate geometric details from global shape to point cloud domain and dual-refinement strategies to iteratively enhance global shape structures and local details; Fine-grained CLIP [song2023fine] exploits textual cues in combination with images for more contextually aware shape completion; DMF-Net [mao2024dmf] and EGIInet [xu2024explicitly] introduce dual-channel modality fusion and shape-aware upsampling to ensure finer geometric consistency for further performance improvement.
@@ III Our Approach
3D Encoder. Taking an incomplete point cloud as input, the initial features are extracted using a 3D encoder. Our 3D encoder follows a modular design, allowing the use of various feature extractors such as PointNet++ [qi2017pointnet++] and DGCNN [wang2019dynamic]. Inspired by SnowflakeNet, our 3D encoder adopts a Set Abstraction (SA) layer with K-NN. Unlike SnowflakeNet, which retains only the final feature, our encoder retains both intermediate feature representations and the final encoded feature, as shown in the orange and green boxes in Figure 1. In contrast to the single-feature extraction in SnowflakeNet and XMFNet, our 3D encoder extracts hierarchical features at multiple levels, capturing both global structural information and local fine-grained details from the input point cloud. In the two-branch configuration, each branch processes the input point cloud independently through the 3D encoder, ensuring diverse feature extraction. These multi-scale representations are then grouped by hierarchical level before being processed in the self-fusion network.
@@
Point Cloud Generation Decoder. The decoder reconstructs the complete point cloud by estimating the missing points and integrating them with the input partial cloud. Since decoding the acquired fusion features requires flexibility and a strong learning capacity, we adopt a decoder architecture similar to XMFNet [wang2022pointattn], designed to process features and learn their implicit shape representations to predict 3D coordinates. Our plug-and-play attention-enhanced self-fusion network is compatible with various other point cloud generation decoders, such as one used by PointAttN [wang2022pointattn].
@@ IV-A Datasets
All experiments are conducted on the ShapeNet-ViPC dataset, which contains 38,328 objects from 13 categories. Each object is associated with 24 partial point clouds and corresponding images, generated under 24 viewpoints, following the previous works’ standard setting [zhang2021view, zhu2023csdn, XMFNet].
@@ IV-B Evaluation Metrics
Following the literature, the performance is evaluated using Chamfer Distance (CD) and F-Score. CD measures the average squared distance between points in the predicted and ground truth point clouds, assessing reconstruction accuracy. F-Score computes overlap precision and recall at a fixed threshold, reflecting structural similarity [zhang2021view, xu2024explicitly].
@@
[…]
@@ IV-C Performance
Table I and Table II summarize our quantitative comparison results for ShapeNet-ViPC dataset, where we report the best officially reported performance of each method based on prior work [xu2024explicitly]. In these comparisons, we evaluate our method alongside existing view-guided point cloud completion approaches, including ViPC [zhang2021view], CSDN [zhu2023csdn], XMFNet [XMFNet] and EGIInet [xu2024explicitly]. Additionally, we benchmark against general point cloud completion methods, such as AtlasNet [groueix2018papier], FoldingNet [yang2017foldingnet], PCN [yuan2018pcn], TopNet [Tchapmi_2019_CVPR], PF-Net [Huang_2020_CVPR], MSN [liu2020morphing], GRNet [xie2020grnet], PoinTr [yu2021pointr], PointAttN [wang2022pointattn], SDT [zhang2022point] and Seedformer [zhou2022seedformer]. Our solution consistently outperforms methods mentioned above in terms of both CD and F-score, demonstrating strong reconstruction accuracy across multiple categories.
@@ IV-E Ablation Study
To assess the impact of different point feature feature extractors and generation decoders on this task, we conduct ablation studies using different combinations of feature extractors (DGCNN [wang2019dynamic] and PointNet++ [qi2017pointnet++]) and generation decoders (XMFNet-based and Transformer-based [wang2022pointattn]). As shown in Table IV, the self-fusion network demonstrates strong compatibility with different feature encoders and point cloud generation decoders, highlighting its flexibility and robustness. This study further validates the generality of the proposed self-fusion network across different architectural configurations.
@@
As extensively studied, vanilla CD is vulnerable to outliers, leading to the development of several CD variants including DCD [wu2021densityaware], HyperCD [lin2023hyperbolic], InfoCD [lin2023infocd] and LandauCD [lin2024loss]. To assess the impact of these variants as loss functions, we conduct an ablation study using a two-branch configuration, with results summarized in Table V. Our findings suggest that, without hyperparameter tuning, these CD variants provide only marginal improvements. Given the substantial effort required for hyperparameter optimization and our primary focus on pipeline design, we adopt vanilla CD as the default loss function for simplicity and efficiency.